A proton with a speed of 2.0x10^5m/s enters a region of space in which source charges have created an electric field potential "v"; What is the proton speed after it move through a potential difference of 100volts.

- What will be the final speed if the proton replace an electron.

Mass of Proton = 1.67x10^-27kg

Mass of Electron = 9.1x10^-30kg


1
Expert's answer
2022-02-28T11:52:11-0500

Given:

"v_i=2.0*10^5\:\rm m/s"

"V=100\:\rm V"


The energy-work theorem says

"\frac{mv_2^2}{2}-\frac{mv_1^2}{2}=eV"

"v_2=\sqrt{v_1^2+2eV/m}"

"v_2=\sqrt{(2.0*10^5)^2+\frac{2*1.6*10^{-19}*100}{1.67*10^{-27}}}"

"v_2=2.43*10^5\:\rm m/s"